# Towards Federated Graph Learning in One-shot Communication

### v5

@@ A.2.1 Performance Comparison
We also compare our method with ReNode (Chen et al., 2021), a representative graph imbalance learning method, which targets to train robust graph models under imbalanced graph data. The comparison results are shown in Table 7. The results show that ReNode cannot effectively model the imbalanced graph data in our scenarios.
@@
A.2.2 Compared with Multi-Rounds FL Methods
To demonstrate the superiority of our method, we further compare our method with multi-round federated learning methods including FedAvg, FedPUB, FedGTA, FedTAD, and FedSpray. We set the communication round of the 5 methods to 100. The experimental results on 5 multi-scale graph datasets under Louvain and Metis partitions are shown in Table 9 and Table 10, respectively. Our method still outperforms in most cases and remains the best average performance.
@@ A.2.2 Compared with Multi-Rounds FL Methods
A.2.3 Full Ablation Study
@@
We conduct the ablation study to evaluate the performance gain of our method. We compare our method with 5 variants: (1) ft: we only fine-tune $M_{G}$ to obtain $M_{i}$ on local graph data, (2) ft_imb: we fine-tune with weighted cross entropy loss to handle data imbalance, (3) ft_dist: in the second stage fine-tuning, we perform distillation on all nodes by setting all $\gamma_{i}$ to a fixed value in Eq. 17; (4) nd_ft_dist: we perform node adaptive distillation along with fine-tuning, without HRE, (5) ours: we perform the complete method (including HRE and personalized training with node-adaptive distillation), and (6) hre_nd_dist: we perform the complete method but without fine-tuning based on $M_{G}$ from the first stage, instead training a model from scratch with cross-entropy loss and node-adaptive distillation.
@@
Figure. 6 and Figure. 8 show our experimental results of accuracy and F1-macro under Louvain and Metis partition, respectively. We make the following 5 observations. (O1) Comparing ft and ft_focal, the focal loss in ft_imb cannot effectively address the non-IID graph data. We believe the reasons are that it’s hard to determine the proper weights under extreme non-IID scenarios. (O2) Comparing ft and ft_dist, a fixed distillation cannot balance the improvements of accuracy and F1-macro. We believe the reason is that it does not distinguish the majority and minority in the distillation. For nodes in major classes with high homophily, the global knowledge of $M_{G}$ may have a negative impact, as demonstrated by the left panel in Figure 5. (O3) Comparing ft and nd_ft_dist, the node adaptive distillation is more flexible in determining when to introduce the global knowledge of $M_{G}$ in fine-tuning. Thus nd_ft_dist achieves better accuracy and F1-macro. (O4) Comparing nd_ft_dist and ours, the HRE strategy further improves performance by generating a higher-quality global pseudo-graph through more precise estimations. (O5) Comparing ours and hre_nd_dist, we conclude that $M_{G}$ provides better and necessary start-point for fine-tuning. Training the local personalized model from scratch, even with node-adaptive distillation and HRE, leads to a significant performance drop.
Figure 8: Ablation study under Metis partition.
@@
We make a further ablation at the client level in Figure 5, evaluating different methods on a client using the CiteSeer dataset with the Louvain partition. As shown in the middle figure, class 3 and class 5 are regarded as major classes due to their larger quantity and relatively high homophily, while class 2 and class 1 are regarded as minor classes with fewer nodes and lower homophily. We make the following 3 observations. (O1) Compared to $M_{G}$, the ft method shows performance gain on major class 3 (+1.57%) and class 5 (+0.56%) but degrades performance on minor class 4 (-1.09%), class 2 (-2.78%), and class 1 (-3.22%). The performance differences after fine-tuning, shown in the left figure, align with the $H$ values in the middle figure, highlighting the limitations of fine-tuning. (O2) The ft_dist method cannot handle this limitation effectively. With a fixed $\gamma_{i}$, the performance on major class 5 worsens (-0.42%), and the performance on minor classes still degrades. This phenomenon motivates us to develop node-adaptive distillation. (O3) Our method, with Class-aware Distillation Factor $w_{dist}$ and node adaptive distillation, boosts performance on all classes by more than +2.4%.
@@
A.2.4 Performance under Model Heterogeneity
we conduct experiments on 10 clients with heterogeneous models. The model configuration is detailed in Section A.1. Under model heterogeneity, most pFGL methods (except FedTAD and FedSpray) do not work since they rely on the model parameter aggregation. Thus we compare our method with FedTAD, FedSpray, DENSE, and Co-Boost. The results under Louvain and Metis partitions are presented in Table 3 and Table 11. The results demonstrate the superior performance of our method in both accuracy and F1-macro.
@@ A.2.4 Performance under Model Heterogeneity
Metis Cora CiteSeer ogbn-arxiv
@@
A.2.5 Inductive Performance
We conduct experiments on three inductive graph datasets: Flickr, Reddit, and Reddit2, and vary the number of clients. Since Flickr has lower homophily, we use 2-layer SIGN (Frasca et al., 2020) for better graph learning. With 10 clients, experimental results under Louvain and Metis partitions are shown in Table 2 and Table 13, respectively. With 20 clients, experimental results under Louvain and Metis partitions are shown in Table 12 and Table 14, respectively. Our method is scalable and generally performs best.
@@ A.2.5 Inductive Performance
Louvain Flickr Reddit Reddit2
@@
Metis Flickr Reddit Reddit2
@@
Metis Flickr Reddit Reddit2
@@ A.2.6 More Datasets
Metis+ Computers Photo CS Physics
@@
A.2.7 Different GNN Architectures
To demonstrate the generality of our proposed method under different graph models, we conduct experiments with additional 3 prevalent graph models including 1) GraphSage (Hamilton et al., 2017), 2) GAT (Veličković et al., 2017), and 3) SGC (Wu et al., 2019). The experimental results under Louvain and Metis partitions are shown in Table 17 and Table 18, respectively. Our method show consistent advantages over other methods.
@@ A.2.7 Different GNN Architectures
A.3 Complexity Analysis
@@
We provide complexity analysis of methods from three perspectives: communication, computation, and memory, with a 2-layer GCN model as an example.
@@
Communication Analysis
We provide communication analysis of methods in Table 19. FedAvg, FedPUB, FedTAD, DENSE, and Co-Boost all need to upload and download the whole model parameters ($O(fd+dc)$, where $d$ is the hidden dimension). FedGTA need to upload additional mixed moments of neighboring features ($O(kKc)$, where $K$ is the order of the moment and $k$ is the propagation steps of Label Propagation) for personalized aggregation. FedSpray needs to upload and download the global feature-structure encoder and structure proxy ($O(fd^{\prime}+d^{\prime}c)$, where $d^{\prime}$ is the output dimension of the encoder and dimension of the structure proxy). Our method only needs to upload the class-wise statistics ($O(chf)$) and download the generated global pseudo-graph ($O(n^{\prime}f+n^{\prime 2})$, $n^{\prime}$ is the number of nodes in the global pseudo-graph, which is usually a small value as indicated in Section A.1). The communication costs of our method are independent of the model size. The larger the model gets, the more communication-efficient our method is compared with other methods.
@@
Computation Analysis
We provide computation analysis of methods in Table 20. we slightly abuse the notation by redefining the number of clients as $N$, the number of edges in local graph data as $m$, and the number of nodes in local graph data as $n$. $K^{\prime}$ is the number of proxy graphs generated in FedPUB and $n^{\prime}$ is the number of nodes in each proxy graph. $p$ is the diffusion step in FedTAD and $O(pn^{2})$ is the complexity to calculate the topology embedding, which could be highly computation-expensive. For FedTAD, DENSE, and Co-Boost, $O(n^{\prime 2}f)$ is the complexity of constructing the generated graph. $L^{\prime}$ is the number of layers of the generator, $f^{\prime}$ is the dimension of the latent features in the generator and $n^{\prime}$ is the number of generated nodes. $O(L^{\prime}n^{\prime}d^{\prime 2})$ is the complexity of forwarding of the generator in the server. For our method, $O(NChf)$ is the complexity of aggregating the statistics. $m^{\prime}$ is the estimated number of edges in the generated global pseudo-graph, usually a small value. $O(m^{\prime}f+Chf)$ is the complexity of calculating alignment loss. $O(n^{\prime 2}f)$ is the complexity of calculating the smoothness loss. The computation complexity on clients in our method is kept as least as possible, the same as the vanilla FedAvg and Standalone.
@@
We provide memory analysis of methods in Table 21. Standalone, FedAvg, FedPUB, DENSE, and Co-Boost only need to store the model parameters ($O(fd+dC)$). FedTAD needs to additionally store the topology embedding calculated by graph diffusion ($O(pn^{2})$), thus it faces OOM on larger graphs. FedSpray needs to store additional encoders and structure proxy. For FedTAD, DENSE, Co-Boost, and our method, the server needs to additionally store the generated graph, which costs approximately $O(n^{\prime}f+n^{\prime 2})$ memory. Since $n^{\prime}$ is usually a small value, it would incur much memory overhead on the server.
A.4 An Equivalent Server-Efficient Variant Method

### v6

@@ A.3.1 Performance Comparison
We also compare our method with ReNode [6], a representative graph imbalance learning method, which aims to train robust graph models under imbalanced graph data. The comparison results are shown in Table 7. The results show that ReNode cannot effectively model the imbalanced graph data in our scenarios.
@@
A.3.2 Compared with Multi-Rounds FL Methods
To demonstrate the superiority of our method, we further compare our method with multi-round federated learning methods, including FedAvg, FedPUB, FedGTA, FedTAD, FedSpray, and FedLoG. We set the communication round of these 6 methods to 100. The experimental results on 5 multi-scale graph datasets under Louvain and Metis partitions are shown in Table 9 and Table 10, respectively. Our method still outperforms in most cases and remains the best average performance.
@@ A.3.2 Compared with Multi-Rounds FL Methods
A.3.3 Detailed Ablation Study
@@
We conduct the ablation study to evaluate the performance gain of our method. We compare our method with 4 variants: (1) ft: only fine-tuning $M_{G}$ to obtain $M_{i}$ on local graph data, (2) ft(focal): fine-tuning with focal loss [39] to handle data imbalance, (3) ft(fixed $\gamma$): in the second stage, performing distillation with a fixed $\gamma$ for all nodes; (4) ours w.o. HRE: performing node adaptive distillation along with fine-tuning, without HRE, and (5) ours: performing the complete method (including HRE).
@@
Figure 4 and Figure 4 show our experimental results of accuracy and F1-macro under Louvain and Metis partition, respectively. We make the following 4 observations. (O1) Comparing ft and ft(focal), the focal loss in ft(focal) cannot effectively address the non-IID graph data. We believe the reasons are that it’s hard to determine the proper weights under extreme non-IID scenarios. (O2) Comparing ft and ft(fixed $\gamma$), a fixed $\gamma$ in distillation cannot balance the improvements of accuracy and F1-macro. We believe the reason is that it does not distinguish the majority and minority in the distillation. For nodes in major classes with high homophily, the global knowledge of $M_{G}$ may have a negative impact, as demonstrated by the left panel in Figure 5. (O3) Comparing ft and ours w.o. HRE, the node adaptive distillation is more flexible in determining when to introduce the global knowledge of $M_{G}$ in fine-tuning. Thus ours w.o. HRE achieves better accuracy and F1-macro. (O4) Comparing ours w.o. HRE and ours, the HRE strategy further improves performance by generating a higher-quality global pseudo-graph through more precise estimations.
@@
We make a further ablation at the client level in Figure 5, evaluating different methods on a client using the CiteSeer dataset with the Louvain partition. As shown in the middle figure, class 3 and class 5 are regarded as major classes due to their larger quantity and relatively high homophily, while class 2 and class 1 are regarded as minor classes with fewer nodes and lower homophily. We make the following 3 observations. (O1) Compared to $M_{G}$, the ft method shows performance gain on major class 3 (+1.57%) and class 5 (+0.56%) but degrades performance on minor class 4 (-1.09%), class 2 (-2.78%), and class 1 (-3.22%). The performance differences after fine-tuning, shown in the left figure, align with the $H$ values in the middle figure, highlighting the limitations of fine-tuning. (O2) The ft(fixed $\gamma$) method cannot handle this limitation effectively. With a fixed $\gamma$, the performance on major class 5 worsens (-0.42%), and the performance on minor classes still degrades. This phenomenon motivates us to develop node-adaptive distillation. (O3) Our method, with Class-aware Distillation Factor $w_{dist}$ and node adaptive distillation, boosts performance on all classes by more than +2.4%.
@@
A.3.4 Performance under Model Heterogeneity
We conduct experiments on 10 clients with heterogeneous models. The model configuration is detailed in Section A.2. Under model heterogeneity, most pFGL methods (except FedTAD and FedSpray) do not work since they rely on the model parameter aggregation. Thus, we compare our method with FedTAD, FedSpray, DENSE, and Co-Boost. The results under Louvain and Metis partitions are presented in Table 3 and Table 11. The results demonstrate the superior performance of our method in both accuracy and F1-macro.
@@ A.3.4 Performance under Model Heterogeneity
Method Cora CiteSeer ogbn-arxiv
@@
Method Flickr Reddit Reddit2
@@
Method Flickr Reddit Reddit2
@@
A.3.5 Inductive Performance
We conduct experiments on three inductive graph datasets: Flickr, Reddit, and Reddit2, and vary the number of clients. Since Flickr has lower homophily, we use the 2-layer SIGN [13] for better graph learning. With 10 clients, experimental results under Louvain and Metis partitions are shown in Table 2 and Table 13, respectively. With 20 clients, experimental results under Louvain and Metis partitions are shown in Table 12 and Table 14, respectively. Our method is scalable and generally performs best.
@@ A.3.5 Inductive Performance
Method Flickr Reddit Reddit2
@@ A.3.6 More Datasets
Method Computers Photo CS Physics
@@
A.3.7 Different GNN Architectures
To demonstrate the generality of our proposed method under different graph models, we conduct experiments with 3 other prevalent graph models, including 1) GraphSage [19], 2) GAT [45], and 3) SGC [49]. The experimental results under Louvain and Metis partitions are shown in Table 17 and Table 18, respectively. Our method shows consistent advantages over other methods.
@@ A.3.7 Different GNN Architectures
A.3.8 Hyper-parameters Analysis
Table 19: Performance with different number of nodes per class
# nodes per class Cora CiteSeer PubMed
Acc. F1 Acc. F1 Acc. F1
1 (used in our paper) 76.43 $\pm$1.24 61.58 $\pm$2.16 71.61 $\pm$0.40 58.24 $\pm$0.96 82.71 $\pm$0.03 66.40 $\pm$0.34
2 70.40 $\pm$2.62 54.29 $\pm$3.45 66.86 $\pm$0.56 53.63 $\pm$1.48 82.57 $\pm$0.04 66.01 $\pm$0.28
3 70.71 $\pm$1.58 52.33 $\pm$2.85 66.39 $\pm$2.73 51.61 $\pm$3.92 82.30 $\pm$0.04 65.17 $\pm$0.30
Table 20: Results on ogbn-arxiv with different number of nodes per class (percent of train set)
# nodes per class (percent of train set) ogbn-arxiv
Acc. F1
0.05% 70.51 $\pm$0.10 34.86 $\pm$0.27
0.1% 70.59 $\pm$0.15 35.28 $\pm$0.27
0.25% (used in our paper) 70.93 $\pm$0.59 36.66 $\pm$0.22
0.5% 70.98 $\pm$0.16 36.67 $\pm$0.25
1% 71.02 $\pm$0.05 36.68 $\pm$0.26
Impact of the pre-set number of nodes in synthesized global pseudo graph.
When increasing the number, the generated global pseudo-graph could have more potential to capture more information. But a large number of nodes would incur optimization difficulties and high communication costs when distributing them to clients. Thus, our principles to choose the number are (1) set the smallest number to capture enough information, (2) keep a similar label distribution with the training set. According to Section A.2, we pre-set the number of nodes to synthesize to 1 for three small graphs and pre-set the number of nodes to synthesize to $p$ percent of the total amount of nodes in corresponding class in the training set, while we keep $p$ as small as possible. We conduct experiments on 10 clients under the Louvain partition to show how the number of nodes impacts the performance. The experimental results are shown in Table 19 and Table 20. For small-scale graphs: Cora, CiteSeer, and PubMed, we pre-set the smallest number of nodes for each class (i.e., one node for each class) and find that it’s already capable. Increasing the number of nodes would not bring benefit since it incurs optimization difficulty. For larger graph: ogbn-arxiv, we pre-set the number of nodes for each class as 0.25% of the total number of corresponding class’s nodes in the train set. Fewer nodes (0.1%, 0.05%) could not fully capture the information in the aggregated class-wise statistics, and lead to a performance drop. More nodes (0.5%, 1%) are more capable but incur more computation and communication costs, and the performance improvement is marginal. Note that the number of nodes in the global pseudo-graph is always small. Even on the largest graph ogbn-product, the total number of nodes of all classes is less than 150. Thus, the memory and communication costs are negligible.
Table 21: Performance with different values of $f_{th}$ and $d_{th}$
ogbn-arxiv $d_{th}=10$ $d_{th}=15$ $d_{th}=20$ $d_{th}=25$
Acc. F1 Acc. F1 Acc. F1 Acc. F1
$f_{th}=0.95$ 70.91 $\pm$0.16 36.21 $\pm$0.32 70.93 $\pm$0.59 36.66 $\pm$0.22 70.94 $\pm$0.15 36.47 $\pm$0.23 70.94 $\pm$0.08 36.37 $\pm$0.06
$f_{th}=1$ 70.82 $\pm$0.07 35.99 $\pm$0.11 70.79 $\pm$0.12 36.14 $\pm$0.51 70.72 $\pm$0.02 35.78 $\pm$0.11 70.84 $\pm$0.22 36.30 $\pm$0.61
Figure 8: Performance with different $K$ values
Table 22: Performance without $f_{th}$ and $d_{th}$
Acc. F1
Ours 76.43 $\pm$1.24 61.58 $\pm$2.16
Without $d_{th}$ (i.e., without (C1)) 75.06 $\pm$0.46 58.75 $\pm$0.82
Without $d_{th}$ and $f_{th}$ (i.e., without (C1) and (C2)) 72.51 $\pm$0.40 53.49 $\pm$2.13
Without $d_{th}$, $f_{th}$ and $topK$ (i.e., without (C1), (C2) and (C3)) 72.48 $\pm$2.45 53.48 $\pm$1.76
Impact of $d_{th}$, $f_{th}$ and $topK$ in HRE.
In our proposed Homophily-guided Reliable node Expansion strategy, we filter out the unreliable soft labels whose node degrees are less than $d_{th}$ (i.e., (C1)) and confidences are less than $f_{th}$ (i.e., (C2)). Also, we only select reliable soft labels that potentially belong to the classes with $topK$ high class homophily (i.e., (C3)). However, our method is not sensitive to the selection of the threshold $f_{th},d_{th}$, and $K$ values. We conduct experiments on ogbn-arxiv dataset under the Louvain partition with 10 clients. The experimental results under different $f_{th}$ and $d_{th}$ are shown in Table 21. We can observe that the performance remains stable under a wide range of $d_{th}$ and both $f_{th}$ values. Regarding $K$ values, each client could select a proper $K$ by statistical analysis of local label distribution, and the performance of our method is not sensitive to $K$ values. The experimental results are in Figure 8 to show the performance stability under different $K$.
To conclude, our method is not sensitive to these hyper-parameters within a proper range in HRE, and the selection would not incur complexity in deployment.
However, these criteria are still necessary because some soft labels produced by label propagation would be unreliable. We provide an experiment in Table 22 on the Cora dataset with 10 clients under the Louvain partition to demonstrate the necessity of $d_{th}$, $f_{th}$, and $topK$.
Impact of $\beta$ in Node Adaptive Distillation
In our proposed Node Adaptive Distillation, we use $\beta$ to control the scale of the value range of the $\gamma_{i}$. We provide experiments on ogbn-arxiv dataset under the Louvain partition with 10 clients to show the performance of our method under different $\beta$ values. The results are in Figure 9. We can observe that although performance would drop under too small or too large $\beta$, our method is not sensitive to $\beta$ as long as its values lie in a proper range.
Figure 9: Performance with different $\beta$ values
A.3.9 Different Adjacency Matrix Implementations
In our method, we derive the adjacency matrix $A^{\prime}$ of the generated global pseudo graph by Eq. 10. We implement the $g_{\theta}$ as a simple 3-layer MLP with input dimensions are $2\times f$ ($f$ is the node feature dimension) and output dimension is 1 (output $\in[0,1]$, represents the probability of existence of a link). Here we compare the other two different implementations of $g_{\theta}$: (1) make $A^{\prime}$ a learnable matrix and directly optimize $A^{\prime}$, (2) $A^{\prime}$ is determined by the dot product of (normalized) node features. The experimental results are shown in Table 23. The advantages of the other two designs cannot be generalized to most datasets. Our design, in contrast, offers flexibility via the learnable MLP and $\delta$ value, thus show overall performance advantages. Note that when setting $\delta=1$, our design fall back to derive a global pseudo graph where all nodes are isolated.
Table 23: Comparison of $g_{\theta}$ implementations
Cora CiteSeer PubMed ogbn-arxiv
Acc. F1 Acc. F1 Acc. F1 Acc. F1
Directly optimize $A^{\prime}$ 70.31 $\pm$4.21 54.36 $\pm$3.71 66.27 $\pm$0.80 53.12 $\pm$0.18 82.71 $\pm$0.16 66.53 $\pm$0.23 70.66 $\pm$0.06 35.54 $\pm$0.44
$A^{\prime}=\langle\mathbf{x}_{i},\mathbf{x}_{j}\rangle$ 73.01 $\pm$1.76 58.78 $\pm$3.01 69.57 $\pm$0.40 58.37 $\pm$0.08 82.68 $\pm$0.05 66.06 $\pm$0.92 70.68 $\pm$0.01 35.30 $\pm$0.54
Ours 76.43 $\pm$1.24 61.58 $\pm$2.16 71.61 $\pm$0.40 58.24 $\pm$0.96 82.71 $\pm$0.03 66.40 $\pm$0.34 70.93 $\pm$0.59 36.66 $\pm$0.22
Table 24: Fine-grained results regarding F1-macro, Recall, and Precision metrics.
total # nodes = 339 class1 (# nodes = 50) class2 (# nodes = 14)
F1-macro Recall Precision F1-macro Recall Precision
Standalone 51.28 50.00 52.63 0 0 0
FedAvg 45.71 40.00 53.33 0 0 0
FedPUB 55.81 60.00 52.17 0 0 0
FedGTA 50.00 55.00 45.83 29.63 57.14 20.00
FedTAD 42.42 35.00 53.85 0 0 0
FedSpray 50.00 56.25 45.00 0 0 0
FedLoG 47.06 40.00 57.14 0 0 0
DENSE 48.65 45.00 52.94 0 0 0
Co-Boost 42.42 35.00 53.85 0 0 0
O-pFGL (Ours) 63.83 75.00 55.56 50.00 42.86 60.00
A.3.10 Fine-grained Experimental Results
In Figure 1, we’ve separately compared the F1-macro of our and baseline methods on minor classes (class 1 and class 2). The class-specific F1 and F1-macro together support that our method not only improves the performance of major classes, but can also effectively generalize to the minority. We select one client among 10 clients on the CiteSeer dataset under the Louvain partition to provide more comparisons on minor classes regarding F1-macro, Recall, and Precision metrics in Table 24.
A.4 Complexity Analysis
@@
We provide the complexity analysis of methods from three perspectives: communication, computation, and memory, with a 2-layer GCN model as an example.
@@
FedLoG $O(fd+dC+n^{\prime 2}f+\text{PG}+\text{Clf}_{h}+\text{Clf}_{t})$ $O(fd+dC+\text{Clf}_{h}+\text{Clf}_{t}+n^{\prime 2}f)$
@@
Communication Analysis.
We provide the communication analysis of the methods in Table 25. FedAvg, FedPUB, FedTAD, DENSE, and Co-Boost all need to upload and download the whole model parameters ($O(fd+dC)$, where $d$ is the hidden dimension). FedGTA needs to upload additional mixed moments of neighboring features ($O(kKC)$, where $K$ is the order of the moment and $k$ is the propagation steps of Label Propagation) for personalized aggregation. FedSpray needs to upload and download the global feature-structure encoder and structure proxy ($O(fd^{\prime}+d^{\prime}C)$, where $d^{\prime}$ is the output dimension of the encoder and dimension of the structure proxy). FedLoG needs to upload the pre-trained prompt generator (for simplicity, we use $O(\text{PG})$ to denote the size of the prompt generator), head-branch classifier and tail-branch classifier (we use $O(\text{Clf}_{h})$ and $O(\text{Clf}_{t})$ to denote the size of the two branch classifiers), synthetic graph ($O(n^{\prime 2}f)$, where $n^{\prime}$ is the pre-set number of nodes), along with the model, to the server. Each client in FedLoG needs to download the aggregated head-branch and tail-branch classifier and the model. The PG, Clfh, and Clft are implemented as multi-layer perceptrons with relatively large hidden dimensions. Thus, FedLoG would incur high upload and download communication costs. In contrast, our method only needs to upload the class-wise statistics ($O(chf)$) and download the generated global pseudo-graph ($O(n^{\prime}f+n^{\prime 2})$). $C$ is the number of classes. $h$ is the propagation step in Eq. 5 and $h\leq 2$ in our experiments. Although in some rare cases where $C$ is extremely large, the upload costs still would not become a huge burden compared with the model size. $n^{\prime}$ is the number of nodes in the global pseudo-graph, which is usually a pre-set small value, as indicated in Sec A.2. The communication costs of our method are independent of the model size. The larger the model, the more communication-efficient our method is compared to other methods.
@@
FedLoG $O(m(f+d)+nd(f+C)+(n^{\prime}+n)d^{\prime 2})$ $O(N(fd+dC+\text{PG}+\text{Clf}_{h}+\text{Clf}_{t}+n^{\prime 2}f))$
@@
Computation Analysis.
We provide computation analysis of methods in Table 26. We slightly abuse the notation by redefining the number of clients as $N$, the number of edges in local graph data as $m$, and the number of nodes in local graph data as $n$. $K^{\prime}$ is the number of proxy graphs generated in FedPUB, and $n^{\prime}$ is the number of nodes in each proxy graph. $p$ is the diffusion step in FedTAD, and $O(pn^{2})$ is the complexity to calculate the topology embedding, which could be highly computationally expensive. For FedTAD, DENSE, and Co-Boost, $O(n^{\prime 2}f)$ is the complexity of constructing the generated graph. $L^{\prime}$ is the number of layers of the generator, $f^{\prime}$ is the dimension of the latent features in the generator, and $n^{\prime}$ is the number of generated nodes. $O(L^{\prime}n^{\prime}d^{\prime 2})$ is the complexity of forwarding the generator in the server. For FedLoG, the additional computation costs come from the prompt generator pretraining and graph synthesis. Although the prompt generator and classifiers are multilayer perceptrons, the pre-training process involves bilevel optimization, which would incur high computation costs in practice. For our method, $O(NChf)$ is the complexity of aggregating the statistics. $m^{\prime}$ is the estimated number of edges in the generated global pseudo-graph, usually a small value. $O(m^{\prime}f+Chf)$ is the complexity of calculating alignment loss. $O(n^{\prime 2}f)$ is the complexity of calculating the smoothness loss. The computation complexity on clients in our method is kept as least as possible, the same as the vanilla FedAvg and Standalone.
@@
FedLoG $O(fd+dC+n^{\prime}f+\text{PG}+\text{Clf}_{h}+\text{Clf}_{t}+n\overline{d}^{k}f)$ $O(N(fd+dC+\text{PG}+\text{Clf}_{h}+\text{Clf}_{t}+n^{\prime 2}f))$
@@
We provide memory analysis of methods in Table 27. Standalone, FedAvg, FedPUB, DENSE, and Co-Boost only need to store the model parameters ($O(fd+dC)$). FedTAD needs to additionally store the topology embedding calculated by graph diffusion ($O(pn^{2})$), thus it faces OOM on larger graphs. FedSpray needs to store additional encoders and structure proxy. FedLoG needs to store the additional prompt generator, two branch classifiers, and generated graphs. Notably, in prompt generator pretraining, FedLoG needs additionally obtain the k-hop subgraph of each node in the local graph, which would incur huge memory ($O(n\overline{d}^{k}f)$, where $\overline{d}$ is the average degree and $k$ is the pre-set number of hops). This is the key factor leading to the out-of-memory issue in dense graphs or large-scale graphs. For FedTAD, DENSE, Co-Boost, and our method, the server needs to additionally store the generated graph, which costs approximately $O(n^{\prime}f+n^{\prime 2})$ memory. Note that $n^{\prime}$ is always a small number. Even on the largest graph ogbn-product, the total number of nodes of all classes is less than 150 ($n^{\prime}<150$). Thus, it would not incur much memory overhead on the server.
A.5 An Equivalent Server-Efficient Variant Method